Domain D = {5, 2, -4}
F (X) = X2 to the second power + 3x + 1
Evaluate the equation of a nonlinear function with two inputs
Please help!!!!! I do not understand

Answer by stanbon(57332)

Domain D = {5, 2, -4}
F(X) = X^2 + 3x + 1
Evaluate the equation of a nonlinear function with two inputs
------------------
F(5) = 5^2 + 3*5 + 1 = 41
------------
F(-4) = (-4)^2 +3(-4) + 1 = 16 -12 +1 = 5
==========================================
Cheers,
Stan H.
===============